QUESTION #22
A SILVERSMITH COMBINES 40 OZ OF PURE SILVER THAT COSTS $8.00 PER OUNCE WITH 200 OZ OF A SILVER ALLOY THAT COSTS $3.50 PER OUNCE. FIND THE SELLING PRICE OF THE MIXTURE.

Answer by checkley77(12844)

40*8+200*3.50=240x
320+700=240x
240x=1,020
x=1020/240
x=$4.25 price of the mix.